# Neuro-Symbolic Video Search

### v1

@@ Abstract
The unprecedented surge in video data production in recent years necessitates efficient tools to extract meaningful frames from videos for downstream tasks. Long-term temporal reasoning is a key desideratum for frame retrieval systems. While state-of-the-art foundation models, like VideoLLaMA and ViCLIP, are proficient in short-term semantic understanding, they surprisingly fail at long-term reasoning across frames. A key reason for this failure is that they intertwine per-frame perception and temporal reasoning into a single deep network. Hence, decoupling but co-designing semantic understanding and temporal reasoning is essential for efficient scene identification. We propose a system that leverages vision-language models for semantic understanding of individual frames but effectively reasons about the long-term evolution of events using state machines and temporal logic (TL) formulae that inherently capture memory. Our TL-based reasoning improves the F1 score of complex event identification by $9-15$% compared to benchmarks that use GPT4 for reasoning on state-of-the-art self-driving datasets such as Waymo and NuScenes. The implementation of our system can be found at Project Webpage.
@@ 1 Introduction
Figure 1: NSVS-TL Pipeline. The input query — “Find the ‘I’m Flying’ scene from Titanic” — is first decomposed into semantically meaningful atomic propositions such as “man hugging woman”, “ship on the sea”, and “kiss” from a high-level user query. SOTA vision and vision-language models are then employed to annotate the existence of these atomic propositions in each video frame. Subsequently, we construct a probabilistic automaton that models the video’s temporal evolution based on the list of per-frame atomic propositions detected in the video. Finally, we evaluate when and where this automaton satisfies the user’s query. We do this by expressing it in a formal specification language that incorporates temporal logic. The TL equivalent of the above query is ALWAYS (normal-$\Box$) “man hugging woman” UNTIL ($\mathsf{U}$) “ship on the sea” UNTIL ($\mathsf{U}$) “kiss”. Formal verification techniques are utilized on the automaton to retrieve scenes that satisfy the TL specification.
There is a significant increase in video data production, with platforms such as YouTube receiving $500$ hours of uploads every minute. Additionally, autonomous vehicle companies such as Waymo generate $10$-$100$ TB of data daily [32], and worldwide security cameras record around $500$ PB each day [31]. Consequently, we require tools with sophisticated query capabilities to navigate this immense volume of video content. For instance, a query such as “Find me all scenes where event A happened, event B did not occur, and event C occurs hours later” requires advanced methods capable of long-term temporal reasoning. Surprisingly, we find that today’s state-of-the-art (SOTA) video and language foundation models fail to identify complex events, especially when key-frames are rare in a video or when the input query is complex (see Fig. 1). Our key insight is that these foundation models aggregate per-frame semantics into a latent vector from which precise scene identification is difficult, especially over long videos. Instead, this paper innovates a novel neuro-symbolic approach that composes neural perception with temporal logic reasoning. We use vision-language models for per-frame semantic understanding but employ state machines and temporal logic to reason about the long-term evolution of events since they inherently capture memory.
Imagine we aim to pinpoint the iconic “I’m flying” scene in the movie Titanic. How do we identify a 5-minute-long scene precisely within the 3-hour and 14-minute movie? Our proposed method, NSVS-TL  presents a solution using a neuro-symbolic approach for video understanding (see Fig. 1). Our key contributions can be summarized as:
@@
    Formal Verification of Video Frame Automaton against a TL Specification: Our algorithm seamlessly provides a confidence measure for predictions on localized scenes by leveraging formal verification on a temporal logic specification. We integrate probabilistic model checkers like STORM [17] to facilitate the verification of the automaton constructed during scene identification and to provide a measure of the confidence of the prediction on the TL specification.
@@
    Datasets: We propose Temporal Logic Video (TLV) datasets, which will be publicly available. These include a synthetic annotated video dataset, created by stitching together images of objects from the ImageNet [7] and COCO [21] datasets, to verify frame search algorithms. Additionally, we annotate publicly available autonomous vehicle datasets – Waymo [32] and NuScenes [13] – with TL specifications derived from detected objects such as “bus”, “pedestrian”, etc., in the scenes.
@@ 2 Related Work
Video Search and Event Detection
Existing research in video event detection predominantly focuses on tracking the spatio-temporal information of objects to identify events, employing deep neural networks for learning latent representations like objects’ motion and position [16, 25, 20, 39, 45, 8, 10, 11]. These methods learn latent representations of videos effectively for downstream event detection or classification, which require substantial computational resources for training neural networks [20, 10, 11]. In contrast, our approach leverages SOTA vision or vision-language models for per-frame semantic recognition, reducing the need for such infrastructure.
Figure 2: Comparative Performance on the Event Identification Task: Video Language Models versus NSVS-TL. The accuracy of event identification with Video Language Models (Blue/Green) drops as video length or query complexity increases. On the other hand, NSVS-TL (Orange) shows consistent performance irrespective of video length or query complexity.
@@
Symbolic Video Understanding
Many works explore approaches to building symbolic representations for video understanding. Such symbolic representations of videos are useful for video classification [12, 36], event detection [25, 39, 20], video question-answering [34, 42], etc. Existing methods either construct graph structures [43, 24, 38] or use latent-space representations as symbolic representations of videos [30, 2, 19]. Among the existing works, several of them propose methods of learning symbolic representation to understand long videos [34, 37, 15, 38]. However, their representations lie in the latent space, which lacks interpretability. In contrast, we use a fully-interpretable automaton-based representation which also provides formal guarantees to downstream tasks like video search.
@@ 3 Preliminaries
Neural Perception Model
@@
  • $F\subseteq Q$ is the set of acceptance or final states representing successful terminations.
In the running example, $Q_{0}$ is defined by the initial detection of “SchoolZoneSign” and “Children”. The accepting states $F$ might contain conditions where the “SchoolZoneSign” is no longer detected, indicating the vehicle’s exit from the school zone. Hence, the PA is a sophisticated model that transitions through states based on the temporal logic specification $\Phi$ (see Eq. 2). The transitions between these states are quantified by $\delta$, with $Q_{0}$ and $F$ framing the start and end of the scenes of interest. $Q$ and $\delta$ are illustrated in Fig. 3 in details. Having the model of video, we now turn to formal verification to ascertain whether this model conforms to our predefined specification $\Phi$.
Formal Verification
@@ 4 Methodology
We introduce a novel way to identify scenes of interest using a neuro-symbolic approach. Given video streams or clips alongside the temporal logic specification $\Phi$, Neuro-Symbolic Visual Search with Temporal Logic (NSVS-TL) identifies scenes of interest through four steps:
  • Step 1: We calibrate the confidence of neural perception models to ensure precise object detection. This calibration enables the detection of relevant propositions in a given frame to construct a probabilistic automaton.
  • Step 2: Subsequently, each frame undergoes a validation process utilizing two distinct validation functions. This step ensures that only frames containing relevant visual information proceed to the next phase of the method.
  • Step 3: Upon validation, we construct a probabilistic automaton dynamically to encapsulate the temporal and logical relations between successive frames.
  • Step 4: Finally, we utilize a model-checking method to determine whether a constructed automaton satisfies the temporal logic specification. If an automaton passes this check, then a sequence of frames within the automaton is identified as a scene of interest by the given temporal logic specification.
@@
We first calibrate neural perception models, $f_{v}$, which detect atomic propositions $p_{k}$ in a frame $\mathcal{F}_{t}$, where $t$ represents the frame’s time index with a confidence $\hat{y}$ for probabilistic verification. We use a generalized logistic function as a mapping estimation function $z(\hat{y})$, which calibrate confidence of models into an accuracy metric, as defined in Eq. 3:
@@
For confidence below $\gamma_{fp}$, it outputs 00, indicating low reliability. Above $\gamma_{tp}$, it outputs $1$, reflecting high accuracy. To determine the accuracy of detection, it utilizes a mapping estimation $z(\hat{y})$ for intermediate confidence levels illustrated in Fig. 4.
@@
Following the calibration step, the process transitions to frame validation. To ensure the presence of atomic propositions, we first validate a frame $\mathcal{F}_{t}$ defined by a detection verification function $V_{c}(\mathcal{F}_{t},p_{k})$:
@@
A frame $\mathcal{F}_{t}$ is considered valid for the automaton if $V_{c}(\mathcal{F}_{t},p_{k})=1$. However, this is not enough to consider the inclusion of $\mathcal{F}_{t}$ in the automaton. For instance, consider a slight modification to our running example:
@@
In this case, we are interested in children walking around the school zone unaccompanied by adults. If we solely rely on the validation function from Eq. 5, frames that include a school sign, children, and adults are added to the automaton. However, including frames with adults contradicts our requirement to exclude them. To address this, we implement symbolic verification to ensure a frame $\mathcal{F}_{t}$ aligns with our temporal logic specification $\Phi$. This process involves evaluating frames against both first-logic operators (denoted as $\Psi$) and temporal operators (denoted as $\Theta$) defined in $\Phi$ as follows:
@@
During a symbolic verification process, each frame initially evaluates the set of first-order logic operators $\Psi$. Formally, $\mathcal{F}_{t}$ is satisfied on $\Psi$ if and only if $\mathcal{F}_{t}\models\Lambda$ as defined below:
  • Conjunction ($\wedge$): A frame $\mathcal{F}_{t}$ is satisfied on $\Psi$ for the conjunction only if both propositions $p_{1}$ and $p_{2}$ associated by $\wedge$ are present in $\mathcal{F}_{t}$: $\mathcal{F}_{t}\models\Lambda$.
  • Disjunction ($\vee$): A frame $\mathcal{F}_{t}$ is satisfied on $\Psi$ for the disjunction if at least one of the propositions $p_{1}$ and $p_{2}$ associated by $\vee$ is present in $\mathcal{F}_{t}$: $\mathcal{F}_{t}\models\Lambda$.
  • Negation ($\neg$): A frame $\mathcal{F}_{t}$ is satisfied on $\Psi$ for the negation if the proposition $p_{k}$ is not present in $\mathcal{F}_{t}$: $\mathcal{F}_{t}\models\Lambda$.
@@
While the first part of symbolic verification, using logical operators ($\Psi$), helps identify frames where children are present without adults ($\neg\mathrm{adult}$), it can ignore the important temporal aspects of our query. Consider a scenario in which our objective is to observe children without adults in the school zone during the period before any adult arrives. If we only validate frames on $\Psi$, a frame meets the immediate logical requirements (children without adults) yet fails to capture the essential temporal aspects — the arrival of adults. This deficiency highlights the importance of integrating the temporal operators ($\Theta$) in our evaluation. Here, we examine the temporal relevance of each proposition in $\Phi$ as follows:
@@
$V_{sv}(\mathcal{F}_{t},P)=\begin{cases}1&\text{if}\hskip 2.84526pt(\exists p_{k}\in P:\tau(p_{k})=1)\vee(\mathcal{F}_{t}\models\Lambda),\\ 0&\text{otherwise}.\end{cases}$ (9)
@@
Once a frame $\mathcal{F}_{t}$ is validated by Eq. 10, we need to build a model of video with validated frames. We call this process Dynamic Automaton Construction because we dynamically construct the automaton based on frame validation instead of building it for an entire video. We first generate $2^{|P|}$ new automaton states, where $|P|$ is the number of atomic propositions. Each of these states correspond to a specific combination of propositions from the set $P$, thereby allowing a comprehensive representation of all possible combinations of propositions within the automaton. In our running example (see Eq. 2), we construct four states representing all possible combinations of atomic propositions or their negation with the proposition set $P=\{p_{1},p_{2}\}$, where $p_{1}=\text{SchoolZoneSign}$ and $p_{2}=\text{Children}$. The labels for these states would be $p_{1}\wedge p_{2},\neg p_{1}\wedge p_{2},p_{1}\wedge\neg p_{2}$, and $\neg p_{1}\wedge\neg p_{2}$, respectively.
@@
TL Specifications Synthetic TLV Dataset Real TLV Dataset
@@
$\diamondsuit$ A
15,750
15,750
$\Box$ A
15,750
15,750
A $\wedge$ B
31,500
A $\mathsf{U}$ B
15,750
15,750
8,736
19,808
(A $\wedge$ B) $\mathsf{U}$ C
31,500
5,789
7,459
Table 1: The number of frames in Synthetic TLV and Real TLV Datasets. It shows the number of video frames generated per TL specification and the source of the data.
5 TLV: Temporal Logic Video Datasets
Given the lack of SOTA video datasets for long-horizon, temporally extended activity and object detection, we introduce the Temporal Logic Video (TLV) datasets. The synthetic TLV datasets are compiled by stitching together static images from image datasets like COCO [21] and ImageNet [7]. This enables the artificial introduction of a wide range of TL specifications. Additionally, we have created two video datasets based on the open-source autonomous vehicle (AV) driving datasets NuScenes [13] and Waymo [32].
5.1 Dataset Compilation
The synthetic TLV datasets are compiled by combining static images randomly selected from the COCO and ImageNet datasets. The propositions, selected from the ground truth labels of original datasets, are used to generate TL specification. For example, if ‘person’ and ‘car’ are selected from the original COCO dataset, they become the propositions of the dataset. Possible TL specification could be either “person and car” or “person until car”. Applying this approach to autonomous driving, we similarly annotate the Waymo and NuScenes datasets, using detected ground truth object labels to create temporal logic annotations. The total number of frames of TLV datasets is detailed in Tab. 1, and key elements of the dataset as follows:
  • The TL specification, expressed in standard temporal logic notation.
  • The list of atomic propositions relevant to the TL formula.
  • The indexed collection of frames that satisfy the specified TL condition.
  • The descriptive labels for each frame, serving as annotations that detail the content or occurrences within the frames.
  • The actual sequence of frames.

### v3

# Towards Neuro-Symbolic Video Understanding

@@ Abstract
The unprecedented surge in video data production in recent years necessitates efficient tools to extract meaningful frames from videos for downstream tasks. Long-term temporal reasoning is a key desideratum for frame retrieval systems. While state-of-the-art foundation models, like VideoLLaMA and ViCLIP, are proficient in short-term semantic understanding, they surprisingly fail at long-term reasoning across frames. A key reason for this failure is that they intertwine per-frame perception and temporal reasoning into a single deep network. Hence, decoupling but co-designing the semantic understanding and temporal reasoning is essential for efficient scene identification. We propose a system that leverages vision-language models for semantic understanding of individual frames and effectively reasons about the long-term evolution of events using state machines and temporal logic (TL) formulae that inherently capture memory. Our TL-based reasoning improves the F1 score of complex event identification by $9-15$%, compared to benchmarks that use GPT-4 for reasoning, on state-of-the-art self-driving datasets such as Waymo and NuScenes. The source code is available at GitHub [^2] https://github.com/UTAustin-SwarmLab/Neuro-Symbolic-Video-Search-Temporal-Logic.
Keywords:
Video Understanding Video Reasoning Neuro Symbolic AI Temporal Logic Formal Methods
@@ 1 Introduction
There is a significant increase in video data production, with platforms such as YouTube receiving $500$ hours of uploads every minute. Additionally, autonomous vehicle companies such as Waymo generate $10$-$100$ TB [28] of data, and worldwide security cameras record around $500$ PB [27] of data daily. Consequently, we require tools with sophisticated query capabilities to navigate this immense volume of video content. For instance, a query such as “Find me all scenes where event A happened, event B did not occur, and event C occurs hours later” requires advanced methods capable of long-term temporal reasoning. Surprisingly, we find that today’s state-of-the-art (SOTA) video and language foundation models fail to identify complex events, especially when key-frames are rare in a video or when the input query is complex (see Fig. 2).
Our key insight is that these foundation models aggregate per-frame semantics into a latent vector from which precise scene identification is difficult, especially over long videos. Instead, this paper innovates a novel neuro-symbolic approach that composes neural perception with temporal logic reasoning. We use vision-language models for per-frame semantic understanding but employ state machines and temporal logic to reason about the long-term evolution of events since they inherently capture memory.
Figure 1: NSVS-TL Pipeline. The input query — “Find the ‘I’m Flying scene from Titanic” — is first decomposed into semantically meaningful atomic propositions such as “man hugging woman”, “ship on the sea”, and “kiss” from a high-level user query. SOTA vision and vision-language models are then employed to annotate the existence of these atomic propositions in each video frame. Subsequently, we construct a probabilistic automaton that models the video’s temporal evolution based on the list of per-frame atomic propositions detected in the video. Finally, we evaluate when and where this automaton satisfies the user’s query. We do this by expressing it in a formal specification language that incorporates temporal logic. The TL equivalent of the above query is ALWAYS ($\Box$) “man hugging woman” UNTIL ($\mathsf{U}$) “ship on the sea” UNTIL ($\mathsf{U}$) “kiss”. Formal verification techniques are utilized on the automaton to retrieve scenes that satisfy the TL specification.
Imagine we aim to pinpoint the iconic “I’m flying” scene in the movie Titanic. How do we identify a 5-minute-long scene precisely within the 3-hour and 14-minute movie? Our proposed method, Neuro-Symbolic Video Search with Temporal Logic (NSVS-TL), presents a solution using a neuro-symbolic approach for video understanding (see Fig. 1). Our key contributions can be summarized as:
@@
    Formal Verification of Video Frame Automaton against a TL Specification: Our algorithm seamlessly provides a confidence measure for predictions on localized scenes by leveraging formal verification on a TL specification. We integrate the probabilistic model checker STORM [13] to formally verify the automaton constructed during scene identification and to measure the probability of satisfying the TL specification.
@@
    Datasets: We propose Temporal Logic Video (TLV) datasets. These include a synthetic annotated video dataset created by stitching together images of objects from the ImageNet [3] and COCO [17] datasets to verify frame search algorithms. Additionally, we annotate autonomous vehicle datasets – Waymo [28] and NuScenes [9] – with TL specifications derived from ground truth object annotations such as “bus”, “pedestrian”, etc., in the scenes.
@@ 2 Related Work
2.1 Video Search and Event Detection
Figure 2: Comparative Performance on the Event Identification Task: Video Language Models versus NSVS-TL. The accuracy of event identification with Video Language Models (Blue/Green) drops as video length or query complexity increases. In contrast, NSVS-TL (Orange) shows consistent performance irrespective of video length or query complexity.
Existing research in video event detection predominantly focuses on tracking the spatio-temporal information of objects to identify events by employing deep neural networks for learning latent representations like objects’ motion and position [12, 21, 16, 34, 41, 4, 8, 7]. These methods learn latent representations of videos for downstream event detection or classification, which require substantial computational resources for training neural networks [16, 8, 7]. In contrast, our approach leverages SOTA vision or vision-language models for per-frame semantic recognition, reducing the need for such infrastructure.
@@ 2.1 Video Search and Event Detection
2.2 Symbolic Video Understanding
Many works explore approaches to building symbolic representations for video understanding. Such symbolic representations of videos are useful for video classification [6, 31], event detection [21, 34, 16], video question-answering [29, 37], etc. Existing methods either construct graph structures [38, 20, 33] or use latent-space representations as symbolic representations of videos [26, 1, 15]. Among the existing works, several of them propose methods of learning symbolic representation to understand long videos [29, 32, 11, 33]. However, these approaches also rely on latent-space representations, which lack interpretability. In contrast, we use a fully interpretable automaton-based representation that provides formal guarantees to downstream tasks like video search.
@@ 3 Preliminaries
3.1 Neural Perception Model
@@ 3.3 Probabilistic Automaton
  • $F\subseteq Q$ is the set of acceptance or terminal states representing successful terminations.
In the running example, $Q_{0}$ is defined by the initial detection of “SchoolZoneSign” and “Children”. The terminal states $F$ might contain conditions where the “SchoolZoneSign” is no longer detected, indicating the vehicle’s exit from the school zone. The PA is a sophisticated model that transitions through states based on the temporal logic specification $\Phi$ (see Eq. 2). The transition probabilities between these states are quantified by $\delta$, with $Q_{0}$ and $F$ denoting the start and end of the scenes of interest. $Q$ and $\delta$ are illustrated in Fig. 3 in details. With this constructed PA, we now conduct formal verification to ascertain whether the PA conforms to our predefined specification $\Phi$.
3.4 Formal Verification
@@ 4 Methodology
We introduce a novel way to identify scenes of interest using a neuro-symbolic approach. Given video streams or clips and a TL specification $\Phi$, Neuro-Symbolic Visual Search with Temporal Logic (NSVS-TL) identifies scenes of interest through four steps:
  • Step 1: We calibrate the confidence of neural perception models to ensure precise object detection. This calibration enables the detection of relevant propositions in a given frame to construct a PA.
  • Step 2: Subsequently, each frame undergoes a validation process with two distinct validation functions. This step ensures that only frames containing relevant visual information proceed to the next phase of the method.
  • Step 3: Upon validation, we construct a probabilistic automaton to encapsulate the temporal and logical relations between successive frames.
  • Step 4: Finally, we utilize model checking to determine whether a constructed automaton satisfies the TL specification. If an automaton passes this check, then a sequence of frames within the automaton is identified as a scene of interest by the given TL specification.
@@
We first calibrate neural perception models, $f_{v}$, which detect atomic propositions $p_{k}$ in a frame $\mathcal{F}_{t}$, where $t$ represents the frame’s time index with a confidence $\hat{y}$ for probabilistic verification. We use a generalized logistic function as a mapping estimation function $z(\hat{y})$, which calibrates the confidence of models into an accuracy metric, as defined in Eq. 3:
@@
For confidence below $\gamma_{fp}$, it outputs 00, indicating low reliability. Above $\gamma_{tp}$, it outputs $1$, reflecting high accuracy. For intermediate confidence levels, it utilizes a mapping estimation $z(\hat{y})$ as illustrated in Fig. 4.
@@
Following the calibration step, the process transitions to frame validation. To ensure the presence of atomic propositions, we first validate a frame $\mathcal{F}_{t}$ defined by a detection verification function:
@@
A frame $\mathcal{F}_{t}$ is considered valid for the automaton if $V_{c}(\mathcal{F}_{t},P)=1$. However, this is not enough to consider the inclusion of $\mathcal{F}_{t}$ in the automaton. For instance, consider a slight modification to our running example:
@@
In this case, we are interested in children walking around the school zone unaccompanied by adults. If we solely rely on the validation function from Eq. 5, frames that include a school sign, children, and adults are added to the automaton. However, including frames with adults contradicts our requirement to exclude them. To address this, we implement symbolic verification to ensure a frame $\mathcal{F}_{t}$ aligns with our TL specification $\Phi$. This process involves evaluating frames against both first-logic operators (denoted as $\Psi$) and temporal operators (denoted as $\Theta$) defined in $\Phi$ as follows:
@@
During a symbolic verification process, each frame initially evaluates the set of first-order logic operators $\Psi$. Formally, $\mathcal{F}_{t}$ is satisfied on $\Psi$, denoted as $\mathcal{F}_{t}\models\Psi$, if and only if the following conditions hold:
  • Conjunction ($\wedge$): A frame $\mathcal{F}_{t}$ is satisfied on $\Psi$ for the conjunction only if both propositions $p_{1}$ and $p_{2}$ associated by $\wedge$ are present in $\mathcal{F}_{t}$.
  • Disjunction ($\vee$): A frame $\mathcal{F}_{t}$ is satisfied on $\Psi$ for the disjunction if at least one of the propositions $p_{1}$ and $p_{2}$ associated by $\vee$ is present in $\mathcal{F}_{t}$.
  • Negation ($\neg$): A frame $\mathcal{F}_{t}$ is satisfied on $\Psi$ for the negation if the proposition $p_{k}$ is not present in $\mathcal{F}_{t}$.
@@
While the first part of symbolic verification, using logical operators ($\Psi$) helps identify frames where children are present without adults ($\neg\mathrm{adult}$), it can ignore the important temporal aspects of our query. Consider a scenario in which our objective is to observe children without adults in the school zone during the period before any adult arrives. If we only validate frames on $\Psi$, a frame meets the immediate logical requirements (children without adults) yet fails to capture the essential temporal aspects — the arrival of adults. This deficiency highlights the importance of integrating the temporal operators ($\Theta$) in our evaluation. Here, we examine the temporal relevance of each proposition in $\Phi$ as follows:
@@
$V_{sv}(\mathcal{F}_{t},P)=\begin{cases}1&\text{if}\hskip 2.84526pt(\exists\hskip 5.69054ptp_{k}\in P:\tau(p_{k})=1)\vee(\mathcal{F}_{t}\models\Psi),\\ 0&\text{otherwise}.\end{cases}$ (9)
@@
Once a frame $\mathcal{F}_{t}$ is validated by Eq. 10, we construct a PA as a model of the video with validated frames. We call this process Dynamic Automaton Construction because we dynamically construct the automaton based on frame validation instead of building it for an entire video. We first generate $2^{|P|}$ new automaton states, where $|P|$ is the number of atomic propositions. Each of these states corresponds to a specific combination of propositions from the set $P$, thereby allowing a comprehensive representation of all possible combinations of propositions within the automaton. In our running example (see Eq. 2), we construct four states representing all possible combinations of atomic propositions or their negation with the proposition set $P=\{p_{1},p_{2}\}$, where $p_{1}=\text{SchoolZoneSign}$ and $p_{2}=\text{Children}$. The labels for these states would be $p_{1}\wedge p_{2},\neg p_{1}\wedge p_{2},p_{1}\wedge\neg p_{2}$, and $\neg p_{1}\wedge\neg p_{2}$, respectively.
@@
5 TLV: Temporal Logic Video Datasets
Given the lack of SOTA video datasets for long-horizon, temporally extended activity and object detection, we introduce the Temporal Logic Video (TLV) datasets. The synthetic TLV datasets are compiled by stitching together static images from image datasets like COCO [17] and ImageNet [3]. This enables the artificial introduction of a wide range of TL specifications. Additionally, we create two video datasets annotated with TL specifications based on the open-source autonomous vehicle (AV) driving datasets NuScenes [9] and Waymo [28]. The dataset is available on GitHub [^3] https://github.com/UTAustin-SwarmLab/Temporal-Logic-Video-Dataset.
5.1 Dataset Compilation
The synthetic TLV datasets are compiled by combining static images randomly selected from the COCO and ImageNet datasets. The propositions selected from the ground truth labels of the original datasets are used to generate TL specifications. For example, if ‘person’ and ‘car’ are selected from the original COCO dataset, they become the propositions for the dataset. Possible TL specifications could be either “person and car” or “person until car”. Applying this approach to autonomous driving, we similarly annotate the Waymo and NuScenes datasets, using detected ground truth object labels to create TL annotations. The total number of frames of TLV datasets is detailed in Tab. 1.
Ground Truth TL Specifications Synthetic TLV Dataset Real TLV Dataset
@@ 5.1 Dataset Compilation
Eventually Event A 15,750 15,750 - -
Always Event A 15,750 15,750 - -
Event A And Event B 31,500 - - -
Event A Until Event B 15,750 15,750 8,736 19,808
(Event A And Event B) Until Event C 31,500 - 5,789 7,459
Table 1: The number of frames in Synthetic TLV and Real TLV Datasets.